OpenStudy (anonymous):

How many terms of the series -12-10-8.... make the sum equal to 48? . . ANSWER: n=16

OpenStudy (anonymous):

Arithmetic Progression/Sequence. \[S _{n}=a+(n-1)d\] \[S _{n}=48\] \[a=-12\] (the first term) \[d=2\] (common difference)

OpenStudy (anonymous):

Find n.

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

i am having problem in simplification

OpenStudy (anonymous):

\[Sn=\frac{ n }{ 2 }[2a+(n-1)d]\]

OpenStudy (anonymous):

i used this formula

OpenStudy (anonymous):

Ah yes sorry about that.

OpenStudy (anonymous):

48=n/2 [2(-12) +(n-1)2]

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

\[48=\frac{ n }{ 2 }[2(-12)+(n-1)2]\]

OpenStudy (anonymous):

Multiply by 2 to both sides.

OpenStudy (anonymous):

When you multiply both side by 2, you get rid of the 2 on RHS(Right Hand Side).

OpenStudy (anonymous):

Expand inside the brackets after that.

OpenStudy (anonymous):

96=n[2(-12)+(n-1)2]

OpenStudy (anonymous):

Yep. Expand the inside first.

OpenStudy (anonymous):

so what did you get inside the brackets. "[..??..]"

OpenStudy (anonymous):

96=n[-24+2n-2]

OpenStudy (anonymous):

Yep Correct. Now expand the outside. So just get rid of the brackets.

OpenStudy (anonymous):

Collect the like terms.

OpenStudy (anonymous):

96=-24n+2n^2-2n

OpenStudy (anonymous):

what will be the quadratic

OpenStudy (anonymous):

Collect the like terms.

OpenStudy (anonymous):

Then move the 96 to the RHS(Right Hand Side).

OpenStudy (anonymous):

Show me final equation of the quadratic.

OpenStudy (anonymous):

Show me the*

OpenStudy (anonymous):

@koli123able

OpenStudy (anonymous):

n^2-13n-48=0

OpenStudy (anonymous):

Good. Now factorise.

OpenStudy (anonymous):

You know the answer is 16

OpenStudy (anonymous):

So then it will have this. (x-16)(x+?)=0

OpenStudy (anonymous):

n^2-16n-3n-48=0

OpenStudy (anonymous):

What are you doing mate?

OpenStudy (anonymous):

Factorise and fill in that question mark. (x-16)(x+?)=0